OpenStudy (anonymous):

Order 34 x 102, 1.2 x 107, 8.11 x 10-3, and 435 from least to greatest.

OpenStudy (anonymous):

here's my answer 1.2x107x 435x 34x102, 8.11x10-3 and the comment my teacher left me Comment: Try converting all into scientific notation to make ordering easier.

OpenStudy (unklerhaukus):

write \[34 \times10^2, 1.2 \times 10^7, 8.11 \times 10^{-3}, 435\] like ``` \[34 \times 10^2, 1.2 \times 10^7, 8.11 \times 10^{-3}, 435\] ``` or like 34 *10^2, 1.2 * 10^7, 8.11 * 10^(-3), 435

OpenStudy (unklerhaukus):

to convert 435 to scientific notation multiply and divide by ten 435=43.5*10=4.35*10^2